# A Fundamental Trade-off in Aligned Language Models and its Relation to Sampling Adaptors

### v2

@@ Abstract
A Fundamental Trade-off in Aligned Language Models and its Relation to Sampling Adaptors
@@ 1 Introduction
This paper seeks to explain the relationship between probability and quality in the specific case of aligned language models, i.e., language models explicitly fine-tuned to better match human preferences, e.g., through reinforcement learning with human feedback (RLHF; Christiano et al., 2017; Leike et al., 2018; Ziegler et al., 2020; Stiennon et al., 2020; Ouyang et al., 2022a, b; Korbak et al., 2022, 2023). In particular, we provide a formal argument and empirical evidence that in RLHF-tuned models, there is an inherent anti-correlation, i.e., a trade-off, between probability and quality. We introduce a new class of sampling methods with useful theoretical properties, globally normalized sampling adaptors, and show that they can be used to control this trade-off by modifying string probability of the generated text.
In the theoretical portion of this paper, we formalize the probability–quality trade-off in aligned language models. Specifically, we show that for corpora of generated strings of a large enough size, the average log-probability under the prior (unaligned) language model ${p}$ trades off with the average score assigned by a reward model (Gao et al., 2023). We refer to this phenomenon as a probability–quality trade-off because, in the context of aligned language models, quality is often operationalized by the reward model (Perez et al., 2022; Lee et al., 2024). This trade-off follows straightforwardly from an application of a concentration inequality and can be seen as a direct corollary of the asymptotic equipartition property (AEP; Cover and Thomas, 2006). In addition to the standard version of the AEP based on Chebyshev’s inequality, we also prove a tighter version of the AEP based on a Chernoff bound that applies to certain language models, including Transformer-based models (Vaswani et al., 2017). Lastly, we show how an appropriate choice of globally normalized sampling adaptor can control the probability of a generated string under the prior ${p}$. The implication of this finding is that they can be used to choose a point on the probability–quality trade-off and control the average quality of generated text. Interestingly, the trade-off predicts the emergence of Simpson’s paradox, which we also observe in our experiments.
In the empirical part of our paper, we present two sets of experiments. First, with synthetic data, we construct toy language and reward models to further demonstrate our theoretical results. This gives easily reproducible empirical evidence for our claim. Then, with a second set of experiments, we show that this trade-off exists in practice with open-sourced RLHF-tuned models, and that globally normalized sampling adaptors do allow us to control where a corpus of generated text will lie on the trade-off. Notably, we find that in practice, locally normalized sampling adaptors (e.g., nucleus sampling; Holtzman et al., 2020; Meister et al., 2023a) are sufficiently close approximations of their globally normalized counterparts, and can be effectively used to control the trade-off.
@@ 2 The Probability–Quality Relationship
Let ${{\Sigma}}$ be an alphabet. A string ${\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}$ is a finite list of symbols drawn from ${{\Sigma}}$. The set of all strings is called ${{\Sigma}}$’s Kleene closure and is denoted as ${{{{\Sigma}}^{*}}}$. A language model ${p}$ is a probability distribution over ${{{{\Sigma}}^{*}}}$. It is often assumed that there is a positive correlation between the quality of a string and its probability under a language model trained on human text ${p}({\boldsymbol{y}})$ (Graves, 2012; Fan et al., 2018; Holtzman et al., 2020; Zhang et al., 2021). [^1] Probability here refers to string likelihood under a general-purpose, unconditional language model. The distinction is important since we will examine the relationship between this probability and the quality of strings in an aligned language model. The idea behind this assumption is that under such a model, a higher probability string should be more human-like, and thus judged by humans to be of higher quality. Because of this positive correlation, string probability is commonly used to reason about text quality in the context of natural language processing. For example, a considerable number of studies on language modeling methods report measures of perplexity to quantify the quality of text produced by the model (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020; Pillutla et al., 2023).
More recently, several authors have observed evidence that this positive correlation breaks down at extremes, finding instead that a string’s probability is positively correlated with its quality up to an inflection point, after which it becomes negatively correlated (Yang et al., 2018; Stahlberg and Byrne, 2019; Holtzman et al., 2020; Zhang et al., 2021; Meister et al., 2022). Meister et al. (2022) show that this inflection point lies near the entropy of human-written text, and hypothesize that a string should encode a similar amount of information to natural language strings to be considered high-quality. This finding has since inspired various sampling adaptors, e.g., locally typical (Meister et al., 2023b) and $\eta$ sampling (Hewitt et al., 2022).
@@ 3 Learning from Human Feedback
Because our investigation focuses on aligned language models, we now introduce RLHF, a popular algorithm for alignment. Given a task of interest, we are interested in producing strings that are aligned with human preferences for that task. Let ${\mathcal{A}}={\{{\textsc{--}},{\textsc{+}}\}}$ – , + } denote binary judgments of alignment, and ${A}$ be an ${\mathcal{A}}$-valued random variable. Then, we assume the existence of a true human-aligned distribution over strings ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{p}({\boldsymbol{y}}\mid{A}={\textsc{+}})$, such that strings with high probability under ${{p}_{{\textsc{+}}}}$ also receive positive scores from human annotators. For example, we may desire that strings that are offensive should have lower probability under ${{p}_{{\textsc{+}}}}$ for chat-related tasks. With this, we can formalize the goal of alignment as obtaining an aligned language model ${q_{{\textsc{+}}}}({\boldsymbol{y}})$ such that ${q_{{\textsc{+}}}}({\boldsymbol{y}})$ is a good approximation to ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})$.
@@ 4 Sampling Adaptors
We are often interested in sampling strings from a language model ${p}({\boldsymbol{y}})$. Sampling is usually performed autoregressively, i.e., we sample a token ${\overline{{y}}}\in{\overline{{{\Sigma}}}}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{{\Sigma}}\cup{\{{\textsc{eos}}\}}$ iteratively from ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$ at each time step $t$ until the special end-of-sequence eos token is reached. Locally normalized sampling adaptors (Meister et al., 2023a) are post-hoc alterations of ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$ that have been shown to dramatically improve the quality of text produced by language models, and are often considered an integral part of a text generation pipeline (Wiher et al., 2022). Common examples include top-${k}$ (Fan et al., 2018) and nucleus sampling (Holtzman et al., 2020).
@@
Applying a locally normalized sampling adaptor to ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$ at each $t$ is commonly formulated as the composition of two steps. First, the application of a transform function ${\boldsymbol{\gamma}}\colon{{\Delta}^{\mid{\overline{{{\Sigma}}}}\mid-1}}\rightarrow{{\mathbbm{R}_{>0}}^{\mid{\overline{{{\Sigma}}}}\mid}}$ that maps the distribution over ${\overline{{{\Sigma}}}}$ to a vector of non-negative values. The transform function is responsible for the core logic of the sampling adaptor, e.g., assigning tokens outside the top-${k}$ zero probability. Second, a normalization step is performed to ensure a valid distribution over ${\overline{{{\Sigma}}}}$, i.e, one where ${\sum_{{\overline{{y}}}\in{\overline{{{\Sigma}}}}}}{\boldsymbol{\gamma}}({p}({\overline{{y}}}\mid{{\boldsymbol{y}}_{<t}}))=1$. When a locally normalized sampling adaptor is applied to a language model ${p}$, we can express the induced distribution ${\ddot{{p}}}$ as
${\ddot{{p}}}({\boldsymbol{y}})$
$=\frac{{\boldsymbol{\gamma}}({p}({\textsc{eos}}\mid{\boldsymbol{y}}))}{{\sum_{{\overline{{y}}}\in{\overline{{{\Sigma}}}}}}{{\boldsymbol{\gamma}}({p}({\overline{{y}}}\mid{\boldsymbol{y}}))}}\frac{{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\boldsymbol{\gamma}}({p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}}))}{{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\sum_{{\overline{{y}}}\in{\overline{{{\Sigma}}}}}}{\boldsymbol{\gamma}}({p}({\overline{{y}}}\mid{{\boldsymbol{y}}_{<t}}))}$ (5)
@@
Despite their empirical success, locally normalized sampling adaptors have several undesirable theoretical properties. First, they can induce a language model that is not tight, i.e., one where probability mass can be placed on infinite-length strings (Welleck et al., 2020; Du et al., 2023). For instance, this can occur if zero probability is placed on eos at every step $t$. Second, the normalization operation at each $t$ can produce unexpected behavior. Specifically, because the denominator used to normalize the local distribution at each $t$ is dependent on the probability mass assigned to other tokens, a string assigned a higher score with the transform function is not necessarily assigned a higher probability under the induced language model ${\ddot{{p}}}$. This can lead to less desirable strings being sampled more frequently. See App. A for an example.
@@
We thus introduce globally normalized sampling adaptors, a procedure to sample from a language model that circumvents these issues. Globally normalized sampling adaptors are also defined with respect to a transform function ${\boldsymbol{\gamma}}$, and thus every existing locally normalized sampling adaptor can be mapped to its globally normalized counterpart. The transform function is similarly applied pointwise at each time step to ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$. But without the normalization step, we define the probability of a string under the induced model ${\widetilde{{p}}}$ as
@@
${{Z}_{\boldsymbol{\gamma}}({{\Sigma}})}$
$={\sum_{{\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}}}{\boldsymbol{\gamma}}\big{(}{p}(\cdot\mid{\boldsymbol{y}})\big{)}({\textsc{eos}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{\boldsymbol{\gamma}}\big{(}{p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\big{)}({{y}_{t}}).$ (7)
@@
That is, the probability of a string under the adapted model is determined only by the transform function, which makes it convenient to reason about how a choice of transform function can modify the properties of the induced distribution ${\widetilde{{p}}}$. Importantly, we use this property in § 5.2 to show that globally normalized sampling adaptors can be used to control the probability–quality trade-off.
To generate strings from a language model using a globally normalized sampling adaptor, we use the Independent Metropolis–Hastings Algorithm (IMHA; Metropolis et al., 1953; Hastings, 1970; Wang, 2022).333 Directly sampling from ${\widetilde{{p}}}$ is not possible because it requires computing the constant ${{Z}_{\boldsymbol{\gamma}}({{\Sigma}})}$. This sum over the countably infinite set of finite strings is in practice unknown. Using IMHA avoids this issue as it only requires the knowledge of the target distribution up to a multiplicative constant. The IMHA is a Markov Chain Monte Carlo (MCMC) method that simulates sampling from a target distribution using a proposal distribution. We detail how the IMHA’s acceptance–rejection protocol can be used to generate text with a globally normalized sampling adaptor in Algorithm 1. The idea is that by sampling sequentially from the proposal (i.e., the distribution induced by a locally normalized sampling adaptor, ${\ddot{{p}}}$) and appropriately accepting or rejecting samples, the generated Markov chain (i.e., the sequence of samples) converges to a stationary distribution equal to the target distribution (i.e., the one induced by a globally normalized sampling adaptor, ${\widetilde{{p}}}$). Convergence is often diagnosed by the absence of autocorrelation between consecutive samples (Deonovic and Smith, 2017). In the case of categorical variables like strings, this can be measured with Cramér’s V (Ialongo, 2016).
@@ 5 Theoretical Results
We are now ready to discuss the theoretical contributions of this paper. In § 5.1 we begin with a formal argument that there exists a fundamental trade-off between the average log-probability under the prior and the average reward for corpora sampled from an aligned language model. Then, in § 5.2, we show how sampling adaptors, by shifting probability mass, we can choose any point on this trade-off. In § 5.3, we conclude the section with a discussion of how this trade-off leads to an emergence of Simpson’s paradox.
@@ 5.1 A Fundamental Trade-off
author=ryan,color=violet!40,size=,fancyline,caption=,]need to check the rest of this proof Prop. 1 follows relatively straightforwardly from Eq. 10 and Eq. 11, when one observes from Eq. 4 that ${q_{{\textsc{+}}}}({\boldsymbol{y}})\propto{p}({\boldsymbol{y}})\exp\left(\frac{1}{\beta}{{r}}({\boldsymbol{y}})\right)$, which implies $\log({q_{{\textsc{+}}}})=\log{p}({\boldsymbol{y}})+\frac{1}{\beta}{{r}}({\boldsymbol{y}})+\mathrm{constant}$. Recall that corpora in the typical set ${T_{N}^{\varepsilon}}({q_{{\textsc{+}}}})$ have average information content $\frac{\log{q_{{\textsc{+}}}}({\boldsymbol{y}})}{N}$ close to the constant pointwise joint entropy ${\mathrm{H}}({\boldsymbol{Y}},{A}={\textsc{+}})$ (Eq. 10). That is, typical corpora by definition exhibit a trade-off between average log-probability under the prior $\frac{\log{p}({\boldsymbol{y}})}{N}$ and average reward $\frac{{{r}}({\boldsymbol{y}})}{N}$. Then, due to Chebyshev’s inequality in Eq. 11, we have $\Big{|}C+\frac{\log{p}({\mathcal{Y}})}{N}+\frac{{{r}}({\mathcal{Y}})}{\beta N}\Big{|}<\varepsilon$ with probability at least $\left(1-\frac{{\mathbb{V}}({I})}{N\varepsilon^{2}}\right)$ for all $N$ and $\varepsilon>0$. When $N\geq\frac{{\mathbb{V}}({I})}{\delta\varepsilon^{2}}$ for some $\delta={\mathcal{O}}(\frac{1}{N})$, the above holds with probability at least $(1-\delta)$, i.e., with high probability, and we arrive at the proposition. The full proof is provided in § B.2. ∎

### v3

# A Probability–Quality Trade-off in Aligned Language Models and its Relation to Sampling Adaptors

@@ Abstract
A Probability–Quality Trade-off in Aligned Language Models and its Relation to Sampling Adaptors
@@ 1 Introduction
This paper seeks to explain the relationship between probability and quality in the specific case of aligned language models, i.e., language models explicitly fine-tuned to better match human preferences, e.g., through reinforcement learning with human feedback (RLHF; Christiano et al., 2017; Leike et al., 2018; Ziegler et al., 2020; Stiennon et al., 2020; Ouyang et al., 2022a, b; Korbak et al., 2022, 2023). In particular, we provide a formal argument and empirical evidence that in RLHF-tuned models, there is an inherent anti-correlation, i.e., a trade-off, between probability and quality.
In the theoretical portion of this paper, we formalize the probability–quality trade-off in aligned language models. Specifically, we show that for corpora of generated strings of a large enough size, the average log-probability under the prior (unaligned) language model ${p}$ trades off with the average score assigned by a reward model (Gao et al., 2023). We refer to this phenomenon as a probability–quality trade-off because, in the context of aligned language models, quality is often operationalized by the reward model (Perez et al., 2022; Lee et al., 2024). This trade-off follows straightforwardly from an application of a concentration inequality and can be seen as a direct corollary of the asymptotic equipartition property (AEP; Cover and Thomas, 2006). In addition to the standard version of the AEP based on Chebyshev’s inequality, we also prove a tighter version of the AEP based on a Chernoff bound that applies to certain language models, including Transformer-based models (Vaswani et al., 2017). Lastly, we show how an appropriate choice of a globally normalized sampling adaptor can control the probability of a generated string under the prior ${p}$. This finding implies that they can be used to choose a point on the probability–quality trade-off and control the average quality of the generated text. Interestingly, the trade-off predicts the emergence of Simpson’s paradox, which we also observe in our experiments.
In the empirical part of our paper, we present two sets of experiments. First, with synthetic data, we construct toy language and reward models to further demonstrate our theoretical results. This gives easily reproducible empirical evidence for our claim. Then, with a second set of experiments, we show that this trade-off exists in practice with open-sourced RLHF-tuned models and that globally normalized sampling adaptors do allow us to control where a corpus of generated text will lie on the trade-off. Notably, we find that in practice, locally normalized sampling adaptors (e.g., nucleus sampling; Holtzman et al., 2020; Meister et al., 2023a) are sufficiently close approximations of their globally normalized counterparts, and can be effectively used to control the trade-off.
@@ 2 The Probability–Quality Relationship
Let ${{\Sigma}}$ be an alphabet. A string ${\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}$ is a finite list of symbols drawn from ${{\Sigma}}$. The set of all strings is called ${{\Sigma}}$’s Kleene closure and is denoted as ${{{{\Sigma}}^{*}}}$. A language model ${p}$ is a probability distribution over ${{{{\Sigma}}^{*}}}$. It is often assumed that there is a positive correlation between the quality of a string and its probability under a language model trained on human text ${p}({\boldsymbol{y}})$ (Graves, 2012; Fan et al., 2018; Holtzman et al., 2020; Zhang et al., 2021). [^1] Probability here refers to string likelihood under a general-purpose, unconditional language model. The distinction is important since we will examine the relationship between this probability and the quality of strings in an aligned language model. The idea behind this assumption is that under such a model, a higher probability string should be more human-like, and thus judged by humans to be of higher quality. Because of this positive correlation, string probability is commonly used to reason about text quality in the context of natural language processing. For example, a number of studies on language modeling methods report measures of perplexity to quantify the quality of text produced by the model (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020; Pillutla et al., 2023).
More recently, several authors have observed evidence that this positive correlation breaks down at extremes, finding instead that a string’s probability is positively correlated with its quality up to an inflection point, after which it becomes negatively correlated (Yang et al., 2018; Stahlberg and Byrne, 2019; Holtzman et al., 2020; Zhang et al., 2021; Meister et al., 2022). Meister et al. (2022) show that this inflection point lies near the entropy of human-written text, and hypothesize that a string should encode a similar amount of information to natural language strings to be considered high-quality. This finding has since inspired various sampling adaptors, e.g., locally typical (Meister et al., 2023b) and $\eta$-sampling (Hewitt et al., 2022).
@@ 3 Learning from Human Feedback
Because our investigation focuses on aligned language models, we now introduce RLHF, a popular algorithm for alignment. Given a task of interest, we are interested in producing strings that are aligned with human preferences for that task. Let ${\mathcal{A}}={\{{\textsc{--}},{\textsc{+}}\}}$ – , + } denote binary judgments of alignment, and ${A}$ be an ${\mathcal{A}}$-valued random variable. Then, we assume the existence of a true human-aligned distribution over strings ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{p}({\boldsymbol{y}}\mid{A}={\textsc{+}})$, such that strings with high probability under ${{p}_{{\textsc{+}}}}$ also receive positive scores from human annotators. For example, we may desire that strings that are offensive should have a lower probability under ${{p}_{{\textsc{+}}}}$ for chat-related tasks. With this, we can formalize the goal of alignment as obtaining an aligned language model ${q_{{\textsc{+}}}}({\boldsymbol{y}})$ such that ${q_{{\textsc{+}}}}({\boldsymbol{y}})$ is a good approximation to ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})$.
@@ 4 Sampling Adaptors
We are often interested in sampling strings from a language model ${p}({\boldsymbol{y}})$. Sampling is usually performed autoregressively, i.e., we sample a symbol ${\overline{{y}}}\in{\overline{{{\Sigma}}}}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{{\Sigma}}\cup{\{{\textsc{eos}}\}}$ iteratively from ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$ at each time step $t$ until the special end-of-sequence eos symbol is reached. Locally normalized sampling adaptors (Meister et al., 2023a) are post-hoc alterations of ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$ that have been shown to dramatically improve the quality of text produced by language models, and are often considered an integral part of a text generation pipeline (Wiher et al., 2022). Common examples include top-${k}$ (Fan et al., 2018) and nucleus sampling (Holtzman et al., 2020).
@@
Applying a locally normalized sampling adaptor to ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$ at each $t$ is commonly formulated as the composition of two steps. First, the application of a transform function ${\boldsymbol{\gamma}}\colon{{\Delta}^{\mid{\overline{{{\Sigma}}}}\mid-1}}\rightarrow{{\mathbbm{R}_{>0}}^{\mid{\overline{{{\Sigma}}}}\mid}}$ that maps the distribution over ${\overline{{{\Sigma}}}}$ to a vector of non-negative values. The transform function is responsible for the core logic of the sampling adaptor, e.g., assigning symbols outside the top-${k}$ zero probability. Second, a normalization step is performed to ensure a valid distribution over ${\overline{{{\Sigma}}}}$, i.e, one where ${\sum_{{\overline{{y}}}\in{\overline{{{\Sigma}}}}}}{\boldsymbol{\gamma}}({p}({\overline{{y}}}\mid{{\boldsymbol{y}}_{<t}}))=1$. When a locally normalized sampling adaptor is applied to a language model ${p}$, we can express the induced distribution ${\ddot{{p}}}$ as
${\ddot{{p}}}($ ${\boldsymbol{y}})=$ (5)
$\frac{{\boldsymbol{\gamma}}({p}({\textsc{eos}}\mid{\boldsymbol{y}}))}{{\sum_{{\overline{{y}}}\in{\overline{{{\Sigma}}}}}}{{\boldsymbol{\gamma}}({p}({\overline{{y}}}\mid{\boldsymbol{y}}))}}\frac{{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\boldsymbol{\gamma}}({p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}}))}{{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\sum_{{\overline{{y}}}\in{\overline{{{\Sigma}}}}}}{\boldsymbol{\gamma}}({p}({\overline{{y}}}\mid{{\boldsymbol{y}}_{<t}}))}$
@@
Despite their empirical success, locally normalized sampling adaptors have several undesirable theoretical properties. First, they can induce a language model that is not tight, i.e., one where probability mass can be placed on infinite-length strings (Welleck et al., 2020; Du et al., 2023). For instance, this can occur if zero probability is placed on eos at every step $t$. Second, the normalization operation at each $t$ can produce unexpected behavior. Specifically, because the denominator used to normalize the local distribution at each $t$ is dependent on the probability mass assigned to other symbols, a string assigned a higher score with the transform function is not necessarily assigned a higher probability under the induced language model ${\ddot{{p}}}$. This can lead to less desirable strings being sampled more frequently. See App. A for an example.
@@
We thus introduce globally normalized sampling adaptors, a procedure to sample from a language model that circumvents these issues. Globally normalized sampling adaptors are also defined with respect to a transform function ${\boldsymbol{\gamma}}$, and thus every existing locally normalized sampling adaptor can be mapped to its globally normalized counterpart. The transform function is similarly applied at each time step to ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$. But without the normalization step, we define the probability of a string under the induced model ${\widetilde{{p}}}$ as
@@
${{Z}_{\boldsymbol{\gamma}}({{\Sigma}})}$ $=$ (7)
${\sum_{{\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}}}{\boldsymbol{\gamma}}\big{(}{p}(\cdot\mid{\boldsymbol{y}})\big{)}({\textsc{eos}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{\boldsymbol{\gamma}}\big{(}{p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\big{)}({{y}_{t}}).$
@@
That is, the probability of a string under the adapted model is determined only by the transform function without additional renormalization, which makes it convenient to reason about how a choice of transform function can modify the properties of the induced distribution ${\widetilde{{p}}}$.
We show in § 5.2 how globally normalized sampling adaptors can be used to control the probability–quality trade-off. To generate strings from a language model using a globally normalized sampling adaptor, we use the Independent Metropolis–Hastings Algorithm (IMHA; Metropolis et al., 1953; Hastings, 1970; Wang, 2022).333 Directly sampling from ${\widetilde{{p}}}$ is not possible because it requires computing the constant ${{Z}_{\boldsymbol{\gamma}}({{\Sigma}})}$. This sum over the countably infinite set of finite strings is in practice unknown. Using IMHA avoids this issue as it only requires the knowledge of the target distribution up to a multiplicative constant. The IMHA is a Markov Chain Monte Carlo (MCMC) method that simulates sampling from a target distribution using a proposal distribution. We detail how the IMHA’s acceptance–rejection protocol can be used to generate text with a globally normalized sampling adaptor in Algorithm 1. The idea is that by sampling sequentially from the proposal (i.e., the distribution induced by a locally normalized sampling adaptor, ${\ddot{{p}}}$) and appropriately accepting or rejecting samples, the generated Markov chain (i.e., the sequence of samples) converges to a stationary distribution equal to the target distribution (i.e., the one induced by a globally normalized sampling adaptor, ${\widetilde{{p}}}$). Convergence is often diagnosed by the absence of autocorrelation between consecutive samples (Deonovic and Smith, 2017). In the case of categorical variables like strings, this can be measured with Cramér’s V (Ialongo, 2016).
@@ 5 Theoretical Results
We are now ready to discuss the theoretical contributions of this paper. In § 5.1 we begin with a formal argument that there exists a fundamental trade-off between the average log-probability under the prior and the average reward for corpora sampled from an aligned language model. Then, in § 5.2, we show how sampling adaptors, by shifting probability mass, can be used to choose any point on this trade-off. In § 5.3, we conclude the section with a discussion of how this trade-off leads to an emergence of Simpson’s paradox.
@@ 5.1 A Fundamental Trade-off
Prop. 1 follows relatively straightforwardly from Eq. 10 and Eq. 11, when one observes from Eq. 4 that ${q_{{\textsc{+}}}}({\boldsymbol{y}})\propto{p}({\boldsymbol{y}})\exp\left(\frac{1}{\beta}{{r}}({\boldsymbol{y}})\right)$, which implies $\log({q_{{\textsc{+}}}})=\log{p}({\boldsymbol{y}})+\frac{1}{\beta}{{r}}({\boldsymbol{y}})+\mathrm{constant}$. Recall that corpora in the typical set ${T_{N}^{\varepsilon}}({q_{{\textsc{+}}}})$ have average information content $\frac{\log{q_{{\textsc{+}}}}({\boldsymbol{y}})}{N}$ close to the constant pointwise joint entropy ${\mathrm{H}}({\boldsymbol{Y}},{A}={\textsc{+}})$ (Eq. 10). That is, typical corpora by definition exhibit a trade-off between average log-probability under the prior $\frac{\log{p}({\boldsymbol{y}})}{N}$ and average reward $\frac{{{r}}({\boldsymbol{y}})}{N}$. Then, due to Chebyshev’s inequality in Eq. 11, we have $\Big{|}C+\frac{\log{p}({\mathcal{Y}})}{N}+\frac{{{r}}({\mathcal{Y}})}{\beta N}\Big{|}<\varepsilon$ with probability at least $\left(1-\frac{{\mathbb{V}}({I})}{N\varepsilon^{2}}\right)$ for all $N$ and $\varepsilon>0$. When $N\geq\frac{{\mathbb{V}}({I})}{\delta\varepsilon^{2}}$ for some $\delta={\mathcal{O}}(\frac{1}{N})$, the above holds with probability at least $1-\delta$, i.e., with high probability, and we arrive at the proposition. The full proof is provided in § B.2. ∎